OpenStudy (anonymous):

If R = (0,0) and T = (2a,0) in equilateral triangle RST what are the coordinates of S in terms of a?

OpenStudy (dumbcow):

distance of RT is: \[d = \sqrt{(2a)^{2} +0^{2}} = 2a\] let x,y be coordinates of S then for RS \[2a = \sqrt{x^{2}+y^{2}}\] for TS \[2a = \sqrt{(x-2a)^{2}+y^{2}}\] Using substitution \[\sqrt{x^{2}+y^{2}} = \sqrt{(x-2a)^{2}+y^{2}}\] \[x^{2} +y^{2} = (x-2a)^{2}+y^{2}\] \[x^{2} = (x-2a)^{2}\] \[x^{2} = x^{2}-4ax +4a^{2}\] \[4ax = 4a^{2}\] \[x = a\] Then \[2a = \sqrt{a^{2}+y^{2}}\] \[4a^{2} = a^{2}+y^{2}\] \[y^{2} = 3a^{2}\] \[y = \sqrt{3}a\] S= (a,sqrt(3)a)